# Language Driven Occupancy Prediction

### v1

@@ Abstract
We introduce LOcc, an effective and generalizable framework for open-vocabulary occupancy (OVO) prediction. Previous approaches typically supervise the networks through coarse voxel-to-text correspondences via image features as intermediates or noisy and sparse correspondences from voxel-based model-view projections. To alleviate the inaccurate supervision, we propose a semantic transitive labeling pipeline to generate dense and fine-grained 3D language occupancy ground truth. Our pipeline presents a feasible way to dig into the valuable semantic information of images, transferring text labels from images to LiDAR point clouds and utimately to voxels, to establish precise voxel-to-text correspondences. By replacing the original prediction head of supervised occupancy models with a geometry head for binary occupancy states and a language head for language features, LOcc effectively uses the generated language ground truth to guide the learning of 3D language volume. Through extensive experiments, we demonstrate that our semantic transitive labeling pipeline can produce more accurate pseudo-labeled ground truth, diminishing labor-intensive human annotations. Additionally, we validate LOcc across various architectures, where all models consistently outperform state-of-the-art zero-shot occupancy prediction approaches on the Occ3D-nuScenes dataset. Notably, even based on the simpler BEVDet model, with an input resolution of $\text{256}\times\text{704}$, LOcc-BEVDet achieves an mIoU of 20.29, surpassing previous approaches that rely on temporal images, higher-resolution inputs, or larger backbone networks. The code for the proposed method is available.
@@ 1 Introduction
Vision-based occupancy prediction aims to estimate both the complete scene geometry and semantics using only image inputs, which serves as a critical foundation for various 3D perception tasks, such as autonomous driving [10, 20], embodied agent [33], mapping and planning [38]. Existing supervised occupancy prediction approaches  [30, 36, 31, 34, 24] are generally constrained to predicting a fixed set of semantic categories, as the dense occupancy ground truth is generated by merging multi-frame LiDAR point clouds with semantic annotations. In this process, human annotators must label each point across thousands of LiDAR frames with a semantic class, which remains highly labor-intensive and costly. Instead, open-vocabulary occupancy (OVO) facilitates predicting the occupancy with arbitrary sets of vocabularies using only unlabeled image-LiDAR data for training [32, 47].
@@
(a) Human-annotated ground truth (b) Voxel-based model-view projection
@@
(c) Image features as intermediates (d) Semantic transitive labeling (ours)
Figure 1: Comparison of Pseudo-Labeled 3D Language Occupancy Ground Truth From Different Pipelines. Human-annotated ground truth requires substantial human effort. Voxel-based model-view projection (b) often introduces labeling errors for occluded voxels, while using image features as intermediates (c) leads to coarse voxel-to-text correspondences. In contrast, our semantic transitive labeling pipeline (d) presents a feasible way to dig into the valuable semantic information of images, transferring texts from images to LiDAR point clouds and ultimately to voxels, producing dense and fine-grained 3D language occupancy ground truth, diminishing laborious human annotations.
Due to the lack of large-scale 3D occupancy datasets with language annotations, distilling knowledge from vision-language models [32, 28, 17, 47, 4] into the occupancy prediction network presents a viable solution. However, the performance of such approaches remains suboptimal, and often demands much computational cost.
To address these issues, we revisit two core aspects for OVO: label transferring process to bridge the gap between language and 3D data, and reconstruction process to assign each voxel a text label. For the label transferring process, using image features from the pretrained vision-language models as intermediates [4, 32] and aligning 3D features with them is a straightforward way to establish indirect correspondences between 3D data and texts. Nevertheless, these vision-language approaches force the vision encoder to align with scene- or region-level texts [28] or produce low-resolution feature maps [40], which lack precise guidance required for pixel-level understanding, resulting coarse voxel-to-text correspondences. Regarding the reconstruction process, previous approaches [32, 47] generally treat all the voxels as surface ones, and project them onto current frames to obtain 3D semantic pseudo labels. But this voxel-based model-view projection overlooks the occlusion problem, leading to noisy and sparse voxel-to-text correspondences.
@@
In this work, we propose a semantic transitive labeling pipeline to generate dense and fine-grained 3D language occupancy ground truth, relieving laborious procedures. In the label transferring process, we use large vision-language models (LVLM)[2, 1, 21] and open-vocabulary segmentation models (OV-Seg)[40, 39, 8] to assign text labels to each pixel, producing high-resolution segmentation maps. Then, through the projection of LiDAR points, the text labels are transferred to the LiDAR points, bridging the gap between 3D data and texts. In the reconstruction process, we model the occlusion relationships of the point clouds during the projection, preventing the wrong label assignment of voxel-based model-view projection. With the pseudo-labeled LiDAR data, we reconstruct a temporally dense scene by merging multi-frame LiDAR points to obtain a dense 3D spatial representation. Furthermore, to reduce the impact of segmentation noise from individual frames, we apply majority-voting voxelization to assign the most frequent label to each voxel. The resulting ground truth enhances 3D language supervision for open-vocabulary occupancy networks, as illustrated in Fig. 1.
@@
Based on the semantic transitive labeling pipeline, we devise LOcc, an effective and generalizable framework that is compatible with most existing supervised occupancy models [12, 11, 20] for OVO. This is accomplished by replacing the original prediction head with a geometry head to predict binary occupancy states and a language head to estimate language features. To efficiently align high-dimensional CLIP [28] embeddings, we devise a language autoencoder that maps them to a low-dimensional latent space, reducing computational cost. We validate our framework on several mainstream approaches, referred to as LOcc-BEVFormer, LOcc-BEVDet, and LOcc-BEVDet4D. All three models consistently surpasses all the previous state-of-the-art zero-shot occupancy prediction approaches on the Occ3D-nuScenes dataset. Notably, even the simpler LOcc-BEVDet, with an input resolution of 256$\times$704, achieves an mIoU of 20.29, outperforming previous methods that rely on temporal image inputs, higher-resolution inputs, or larger backbone networks. In summary, our contributions are summarized as follows:
@@ 2.1 Vision-based Occupancy Prediction
Occupancy prediction jointly estimates the occupancy state and semantic label for every voxel in the scene using images. Voxels are classified as free, occupied, or unobserved, with occupied voxels assigned corresponding semantic labels. Early works [3, 6, 46, 19, 48, 43] typically use monocular or stereo images as inputs, also referred to as semantic scene completion (SSC). With advancements in 3D perception [20, 12, 11], the occupancy of surrounding scenes using multi-camera images [37, 23, 13, 42] has garnered significant attention due to its strong representation capabilities in 3D space. SurroundOcc [36] introduces a new pipeline for constructing dense occupancy ground truth. Occ3D [30] establishes benchmarks for surround occupancy prediction through a three-stage data generation pipeline: voxel densification, occlusion reasoning, and image-guided voxel refinement. OpenOccupancy [34] provides a benchmark with higher-resolution ground truth, while OccNet [31] offers flow annotations.
Although the above approaches provide an automatic data generation pipeline to construct dense occupancy ground truth, they require LiDAR data with semantic annotations. To achieve this, human annotators must label each point across thousands of LiDAR frames with a semantic class, which remains highly labor-intensive and costly. Besides, the valuable semantic information of images are ignored during the occupancy generation process. Instead, in this work, we propose a semantic transitive labeling pipeline to dig into the valuable semantic information from images, transferring text labels from images to LiDAR point clouds and ultimately to voxels. It is compact and does not require labor-intensive human annotations, offering a more cost-effective way to incorporate semantic information from images.
@@ 2.2 Open-Vocabulary Understanding
The recent advances of vision-language models [28] have showed a remarkable zero-shot performance by using only paired images and captions for model training, demonstrating the strong connection between images and natural language. Then, numerous works explore the potential of language driven downstream applications, such as object detection [44, 7], tracking [18], and segmentation [17, 8, 40, 39]. Taking the image features of these 2D foundation models as intermediaries, OpenScene [25] distills 2D vision-language knowledge into a 3D point cloud network and achieve zero-shot 3D segmentation. LERF [15] is the first to embed CLIP features into NeRF for controllable rendering. LangSplat [27] uses SAM [16] to tackle the point ambiguity issue for 3D language field modeling.
@@ 2.3 Open-Vocabulary Occupancy Prediction
Benefiting from the development of vision-language foundation models, OVO enables the prediction of occupancy associated with arbitrary vocabulary sets without requiring manual annotations of extensive 3D voxel data. POP-3D [32] employs only unlabeled images and sparse LiDAR point clouds for training. It distills vision-language knowledge from the pretrained LSeg [17] into the occupancy networks by computing cosine similarity between 3D features and their corresponding 2D features, obtained by projecting the point cloud onto the image plane. VEON [47] introduces a side adaptor network to integrate CLIP [28] into the occupancy prediction model. However, the feature maps from existing vision-language models lack the fine-grained guidance required for pixel-level understanding, failing to establish precise correspondences between voxels and texts. Furthermore, current works typically treat all the voxels as surface ones and employ voxel-based model-view projection to obtain supervision, leading to a lots of misprojections.
@@ 3.1 Pseudo-Labeled Ground Truth Generation
Scene Reconstruction. The pseudo-labeled LiDAR data is then used to perform scene reconstruction to establish correspondences between the texts and the voxels. To combine the multi-frame point clouds, we first transform their coordinates into the world coordinate system using the calibration matrices and ego poses. For the $k$-th frame, we transform the unified point cloud to this frame and apply majority-voting voxelization on it to obtain the pseudo-labeled 3D language occupancy ground truth. Each voxel takes the most frequent text within it as its text label. Compare to the voxel-based model-view projection, our method obtains text label by projecting LiDAR point clouds onto image planes, which are easily calibrated with the corresponding images during the data collection process. Besides, the majority-voting is more robust and less susceptible to the influence of single-point projection errors, generating ground truth with higher metrics compared to the nearest-point voxelization, as shown in the ablation experiments.
@@ 3.2 Open Vocabulary Occupancy Prediction
Image Encoder. The image encoder includes an image backbone network for extracting multi-scale features, complemented by an FPN network to enhance features across different levels. The enhanced features $\mathbf{F}\in{\mathbb{R}^{H\times W\times C}}$ are then used to construct 3D features.
@@
Table 1: 3D occupancy prediction performance on the Occ3D-nuScenes dataset. We report the mean IoU (mIoU) for semantics across different categories, along with per-class semantic IoUs. represents the reproduced results in [42]. The methods marked with $\dagger$ indicate that they specifically design the set of vocabularies for training. The top three results among the methods supervised by pseudo-labeled ground truth are in red, green, and blue, respectively.
@@
Supervised by pseudo-labeled ground truth.
@@
3.3 Language Autoencoder
@@ 4.1 Datasets and Metrics
Datasets. We evaluate LOcc on the nuScenes dataset [5, 30], a large-scale autonomous driving dataset collected across various cities and weather conditions. Each frame includes a LiDAR scan and six RGB images from different viewpoints. It contains 700 training scenes and 150 validation scenes. The spatial volume ranges from $-\text{40}m$ to $\text{40}m$ for the X and Y axes, and $-\text{1}m$ to $\text{5.4}m$ for the Z axis. Voxelization of this volume produces a set of 3D grids with a resolution of $\text{200}\times\text{200}\times\text{16}$, where each voxel measures $\text{0.4}m\times\text{0.4}m\times\text{0.4}m$. The ground truth includes 17 unique labels (16 semantic classes and 1 free class).
@@
4.2 Implementation Details
Network Structures. To demonstrate the generalizability of the proposed LOcc, we evaluate it on BEVFormer [20], BEVDet [12], and BEVDet4D [11], including both forward- and backward-projection methods. For BEVDet [12] and BEVDet4D [11], we employ ResNet-50 [9] as the image backbone. The resolution of the voxel features after the 2D-to-3D transformation is $\text{200}\times\text{200}\times\text{16}$, with a feature dimension of 64. For BEVFormer [20], we adopt ResNet101-DCN [9] as the backbone, with the default size for BEV queries set to $\text{200}\times\text{200}$, with a feature dimension of 256. We remove the temporal self-attention layer and use 6 cross-attention layers. For each query, it corresponds to 4 target points with different heights in 3D space, with height anchors uniformly predefined from $-\text{1}m$ to $\text{5.4}m$. The number of sampling points around each reference point is set to 8. For all models, the image size is resized to either $\text{256}\times\text{704}$ or $\text{512}\times\text{1408}$, and the language features are compressed to a dimension of 128.
Training Setup. We implement the networks using PyTorch. The models are trained for 25 epochs on 8 NVIDIA 3090 GPUs, with a batch size of 1 on each GPU. We employ the AdamW [22] optimizer with $\beta_{1}=\text{0.9},\beta_{2}=\text{0.99}$. The maximum learning rate is set to $\text{3}\times\text{10}^{-\text{4}}$, and the cosine annealing learning rate strategy is adopted for the learning rate decay, where the cosine warm up strategy is applied for the first $\text{5}\%$ iterations.
Table 2: Ablation study on the framework for generating dense and fine-grained pseudo-labeled 3D language occupancy ground truth. For label transferring process, we analyze three settings: (a) using image features from vision-language foundation models as intermediates, (b) using single-frame vocabularies, and (c) integrating vocabularies across multiple consecutive frames. For scene reconstruction process, we replace the majority-voting voxelization with (d) nearest-point voxelization and (e) voxel-based model-view projection. The best results for each model are in red.
Image Feat. Single-frame vocab. Consecutive-frame vocab. Point-based Projection Voxel-based Projection LOcc-BEVFormer LOcc-BEVDet LOcc-BEVDet4D
Voting voxelization Nearest voxelization
(a) 15.04 17.56 19.37
(b) 17.87 19.61 22.15
(c) 18.62 20.29 22.95
(d) 18.32 19.98 22.68
(e) 13.77 14.49 15.41
Table 3: Detailed metrics of the pseudo-labeled ground truth from different generation settings on the Occ3D-nuScenes dataset. The settings match those outlined in Table 2. We report the mean IoU (mIoU) for semantics across different categories. The top three results different settings are in red, green, and blue, respectively.

### v3

@@ Abstract
We introduce LOcc, an effective and generalizable framework for open-vocabulary occupancy (OVO) prediction. Previous approaches typically supervise the networks through coarse voxel-to-text correspondences via image features as intermediates or noisy and sparse correspondences from voxel-based model-view projections. To alleviate the inaccurate supervision, we propose a semantic transitive labeling pipeline to generate dense and fine-grained 3D language occupancy ground truth. Our pipeline presents a feasible way to dig into the valuable semantic information of images, transferring text labels from images to LiDAR point clouds and ultimately to voxels, to establish precise voxel-to-text correspondences. By replacing the original prediction head of supervised occupancy models with a geometry head for binary occupancy states and a language head for language features, LOcc effectively uses the generated language ground truth to guide the learning of 3D language volume. Through extensive experiments, we demonstrate that our transitive semantic labeling pipeline can produce more accurate pseudo-labeled ground truth, diminishing labor-intensive human annotations. Additionally, we validate LOcc across various architectures, where all models consistently outperform state-of-the-art zero-shot occupancy prediction approaches on the Occ3D-nuScenes dataset.
[…]
@@ 1 Introduction
Vision-based occupancy prediction aims to estimate both the complete scene geometry and semantics using only image inputs, which serves as a critical foundation for various 3D perception tasks, such as autonomous driving [10], embodied agent [36, 15], mapping and planning [42, 34]. Existing supervised occupancy prediction approaches [33, 39, 34, 37, 27, 26] are generally constrained to predicting a fixed set of semantic categories, as the dense occupancy ground truth is generated by merging multi-frame LiDAR point clouds with semantic annotations. In this process, human annotators must label each point across thousands of LiDAR frames with a semantic class, which remains highly labor-intensive and costly. Instead, open-vocabulary occupancy (OVO) facilitates predicting the occupancy with arbitrary sets of vocabularies using only unlabeled image-LiDAR data for training [35, 56].
@@
(a) Human-annotated ground truth (b) Image features as intermediates
@@
(c) Voxel-based model-view projection (d) Semantic transitive labeling (ours)
Figure 1: Comparison of pseudo-Labeled 3D language occupancy ground truth from different pipelines.
Due to the lack of large-scale occupancy datasets with language annotations, distilling knowledge from vision-language models [35, 31, 18, 4] into occupancy prediction networks presents a viable solution. Some approaches [35, 18] use image features from the pretrained vision-language models [31, 18] as intermediates [4, 35], subsequently bridging the gap between 3D voxel and language features. However, the feature values of objects with the same semantics may vary, as image features inherently encode both semantic and appearance information, leading to inconsistencies in semantic representations across different inputs. Consequently, the pseudo-labeled ground truth derived from these features is inherently noisy in terms of semantics, failing to provide precise guidance for voxel-level understanding. Moreover, the single-scan sparse LiDAR [35] can only provide sparse geometry supervision, further resulting in sparse and coarse voxel-to-text correspondences. To address this, VEON [56] enhances geometry supervision by employing dense binary occupancy maps, which are generated through offline post-processing. However, it treats voxels as point clouds and assigns semantic labels to them by model-view projection, which overlooks the occlusion problem and only uses single-frame images, leading to coarse voxel-to-text correspondences, too.
@@
To address the aforementioned issues, we revisit the core aspect of OVO: generating dense and fine-grained 3D language occupancy ground truth. In this work, we propose a semantic transitive labeling pipeline to achieve this goal, without laborious procedures. It can be roughly divided into two steps: label transferring and scene reconstruction. For the first step, we use large vision-language models (LVLM) [2, 1, 24] and open-vocabulary segmentation models (OV-Seg) [44, 43, 8] to assign text labels to each pixel, obtaining consistent semantic representations. Then, through the projection of LiDAR points, the text labels are transferred to the LiDAR points, bridging the gap between 3D data and texts. In the reconstruction process, we model the occlusion relationships of the point clouds during point projection, preventing the wrong label assignment of voxel-based model-view projection. With the pseudo-labeled LiDAR data, we reconstruct a temporally dense scene by merging multi-frame LiDAR points to obtain a dense 3D spatial representation. Furthermore, to reduce the impact of segmentation noise from individual frames, we apply majority-voting voxelization to assign the most frequent label to each voxel. The resulting ground truth enhances 3D language supervision for open-vocabulary occupancy networks. A comparison of different pseudo labeled ground truth from different pipelines is illustrated in Fig. 1.
@@
Based on the semantic transitive labeling pipeline, we devise LOcc, an effective and generalizable framework that is compatible with most existing supervised occupancy models [12, 11, 23] for OVO. To efficiently align high-dimensional CLIP [31] embeddings, we devise a language autoencoder that maps them to a low-dimensional latent space, reducing computational cost. We validate our framework on several mainstream approaches, referred to as LOcc-BEVFormer, LOcc-BEVDet, and LOcc-BEVDet4D. All three models consistently surpass all the previous state-of-the-art zero-shot occupancy prediction approaches on the Occ3D-nuScenes dataset. Notably, even the simpler LOcc-BEVDet, with an input resolution of 256$\times$704, achieves a mIoU of 20.29, outperforming previous methods that rely on temporal image inputs, higher-resolution inputs, or larger backbone networks. In summary, our contributions are summarized as follows:
@@ 2.1 Vision-based Occupancy Prediction
Occupancy prediction jointly estimates the occupancy state and semantic label for every voxel in the scene using images. Voxels are classified as free, occupied, or unobserved, with occupied voxels assigned corresponding semantic labels. Early works [3, 6, 55, 22, 57, 52] typically use monocular or stereo images as inputs, also referred to as semantic scene completion (SSC). With advancements in 3D perception [23, 29, 47, 12, 11, 13, 19, 46, 21, 50, 51], the occupancy of surrounding scenes using multi-camera images [40, 41, 26, 13, 49] has garnered significant attention due to its strong representation capabilities in 3D space. SurroundOcc [39] introduces a new pipeline for constructing dense occupancy ground truth. Occ3D [33] establishes benchmarks for surround occupancy prediction through a three-stage data generation pipeline: voxel densification, occlusion reasoning, and image-guided voxel refinement. OpenOccupancy [37] provides a benchmark with higher-resolution ground truth, while OccNet [34] offers flow annotations.
@@ 2.2 Open-Vocabulary Understanding
The recent advances in vision-language models [31] have shown a remarkable zero-shot performance by using only paired images and captions for model training, demonstrating the strong connection between images and natural language. Then, numerous works explore the potential of language driven downstream applications, such as object detection [53, 7], tracking [20], and segmentation [18, 8, 44, 43]. Taking the image features of these 2D foundation models as intermediaries, OpenScene [28] distills 2D vision-language knowledge into a 3D point cloud network and achieves zero-shot 3D segmentation. LERF [16] is the first to embed CLIP features into NeRF for controllable rendering. LangSplat [30] uses SAM [17] to tackle the point ambiguity issue for 3D language field modeling.
@@ 2.3 Open-Vocabulary Occupancy Prediction
Benefiting from the development of vision-language foundation models, OVO enables the prediction of occupancy associated with arbitrary vocabulary sets without requiring manual annotations of extensive 3D voxel data. POP-3D [35] employs only unlabeled images and sparse LiDAR point clouds for training. It distills vision-language knowledge from the pretrained LSeg [18] into the occupancy networks by computing cosine similarity between 3D features and their corresponding 2D features, obtained by projecting the point cloud onto the image plane. VEON [56] introduces a side adaptor network to integrate CLIP [31] into the occupancy prediction model. However, the extracted image features may exhibit inconsistencies across different input images, lacking accurate guidance for voxel-level understanding and leading to coarse voxel-to-text correspondences. Furthermore, current works typically treat the voxels as point clouds and employ voxel-based model-view projection to obtain supervision, leading to a lots of misprojections.
@@ 3.1 Pseudo-Labeled Ground Truth Generation
Scene Reconstruction. The pseudo-labeled LiDAR data is then used to perform scene reconstruction to establish correspondences between the texts and the voxels. To combine the multi-frame point clouds, we first transform their coordinates into the world coordinate system using the calibration matrices and ego poses. In alignment with VEON [56], we distinguish moving and static objects by leveraging the geometric information of 3D bounding boxes. For the $k$-th frame, we transform the unified point cloud to this frame and apply majority-voting voxelization on it to obtain the pseudo-labeled 3D language occupancy ground truth. Each voxel takes the most frequent text within it as its text label, and voxels without point clouds are set to empty. Compare to the voxel-based model-view projection, our method obtains text label by projecting LiDAR point clouds onto image planes, which are easily calibrated with the corresponding images during the data collection process. Besides, the majority-voting is more robust and less susceptible to the influence of single-point projection errors, generating ground truth with higher metrics compared to the nearest-point voxelization, as shown in the ablation experiments.
@@ 3.2 Open Vocabulary Occupancy Prediction
Image Encoder. The image encoder includes an image backbone network for extracting multi-scale features, complemented by an FPN [45] network to enhance features across different levels. The enhanced features $\mathbf{F}\in{\mathbb{R}^{H\times W\times C}}$ are then used to construct 3D features.
@@
3.3 Language Autoencoder
Table 1: Zero-shot occupancy prediction results on the Occ3D-nuScenes dataset. We report the mean IoU (mIoU) for semantics across different categories, along with per-class semantic IoUs. The symbol indicates reproduced results from [49]. The methods marked with $\dagger$ indicate that they specifically design the set of vocabularies for training. The methods marked with $\ddagger$ are implemented the official open-source codebase. The top three results among the methods supervised by pseudo-labeled ground truth are in red, green, and blue, respectively.
@@ 3.3 Language Autoencoder
Supervised by images.
@@
POP-3D-BEVFormer [35] ResNet-101 256 $\times$ 704 11.24 0.00 7.32 0.95 11.23 9.18 2.59 2.44 2.21 0.78 1.7 8.26 56.02 0.00 19.28 21.47 21.54 26.18
POP-3D-BEVFormer [35] ResNet-101 512 $\times$ 1408 12.68 0.00 9.43 2.43 15.21 10.39 3.54 3.78 3.98 2.91 1.96 10.01 57.28 0.00 21.01 23.07 22.67 27.87
POP-3D-BEVDet [35] ResNet-50 256 $\times$ 704 12.30 0.00 8.72 1.94 13.11 11.48 3.12 3.29 3.47 1.78 1.21 9.43 56.81 0.00 20.17 23.29 23.33 28.12
POP-3D-BEVDet [35] ResNet-50 512 $\times$ 1408 12.85 0.00 7.43 2.21 15.13 12.25 3.40 3.47 4.21 2.22 2.42 10.67 58.23 0.00 21.46 24.33 22.16 28.88
POP-3D-BEVDet4D [35] ResNet-50 256 $\times$ 704 14.57 0.00 9.94 3.09 15.83 13.51 4.38 3.57 4.64 3.84 2.21 12.20 59.71 0.00 26.60 28.16 27.88 32.20
POP-3D-BEVDet4D [35] ResNet-50 512 $\times$ 1408 15.64 0.00 10.06 3.63 16.74 14.32 5.38 4.01 5.14 3.52 2.46 15.20 61.13 0.00 28.82 30.16 31.92 33.31
@@
Table 2: Ablation study on the framework for generating dense and fine-grained pseudo-labeled 3D language occupancy ground truth. For label transferring process, we analyze three settings: (a) using image features from vision-language foundation models as intermediates, (b) using single-frame vocabularies, and (c) integrating vocabularies across multiple consecutive frames. For scene reconstruction process, we replace the majority-voting voxelization with (d) nearest-point voxelization and (e) voxel-based model-view projection. The best results for each model are in bold.
Image Feat. Single-frame vocab. Consecutive-frame vocab. Point-based Projection Voxel-based Projection LOcc-BEVFormer LOcc-BEVDet LOcc-BEVDet4D Pseudo-labeled GT
Voting voxelization Nearest voxelization
(a) 15.04 17.56 19.37 22.12
(b) 17.87 19.61 22.15 22.66
(c) 18.62 20.29 22.95 25.53
(d) 18.32 19.98 22.68 23.80
(e) 13.77 14.49 15.41 19.55
@@ 4.1 Datasets and Metrics
[…]
@@
4.2 Evaluation Benchmarks
The evaluation benchmarks focus on two settings, i.e., zero-shot and open-vocabulary. In the zero-shot setting, none of the ground truth classes are seen during training. In contrast, the open-vocabulary setting assumes that only a subset of classes is available during training, and the model is expected to recognize both seen and unseen classes during evaluation
Table 3: Detailed metrics of the pseudo-labeled ground truth from different generation settings on the Occ3D-nuScenes dataset. The settings match those outlined in Table 2. We report the mean IoU (mIoU) for semantics across different categories. The best results of different settings are in bold.